# Operator Growth in Disordered Spin Chains: Indications for the Absence of Many-Body Localization

### v3

@@ III.2.1 No disorder
From the exact solution of the model, we know for sure that the eigenstates of the XXX model are not localized. We can calculate the number of distinct terms $n(k)$ in the commutator $[H,\sigma_{0}^{z}]^{(k)}$ at order $k$ by symbolic calculations up to order $k=18$ exactly. We have not been able to identify the series but asymptotically the number of terms $n(k)$ does grow exponentially with a multiplicative algebraic correction which is consistent with the general bound (II.15) which we derived earlier. The data and a corresponding fit are shown in Fig. 9.
@@
Figure 9: In the interacting model with $\Delta=1$, the number of different terms $n(k)$ in the commutator $[H,\sigma_{0}^{z}]^{(k)}$ grows exponentially (symbols) and is very well fitted by $n(k)\sim 1.514\,k^{0.803}\exp(0.922k)$ in the non-disordered case and $n(k)\sim 1.492\,k^{0.815}\exp(0.998k)$ in the case with disorder (solid lines). The general bounds (II.15) with $f=3$ and $f=4$ are shown as well.
@@
The fit works extremely well with a least square deviation of $\chi^{2}\sim 10^{-2}$. The local operator basis for the Heisenberg model has dimension $f=4$ and consists of the operators ${\mathbbm{1},\sigma^{x},\sigma^{y},\sigma^{z}}$. We note, however, that only a small subset of terms in $[H,\sigma^{z}_{0}]^{(k)}$ contains an identity in between Pauli matrices. Ignoring such terms, we can set $f=3$ in Eq. (II.15) which gives a much tighter bound, see Fig. 9. The exponential growth of $n(k)$, instead of the quadratic growth in the free-fermion case, indicates that in the corresponding commutator graph the number of nodes per level is no longer constant.
@@
We note that the number of distinct terms is different and always smaller than the number of connected clusters which we have calculated in Eq. (II.11). The reason is that different connected clusters can result in the same term. For the calculation of the norm this does not matter though. We can bound the contribution of each connected cluster as in Eq. (II.12) even if some of the resulting terms are the same. In Fig. 10 we show that this general bound indeed describes the norm growth well if we use the bound on the local Hamiltonian, $||h^{a}_{I}||\leq J$, as a fitting parameter.
@@
Figure 10: The 1-norm in the interacting case with $\Delta=1$ grows faster than exponential for all disorder strengths shown, consistent with the rescaled general bound, Eq. (II.12). The data are averaged over $30$ disorder realizations for $D=2.5$, $D=5$ and $D=12.5$ and over 50 realizations for $D=25$.
@@
Figure 12: The norm of operators $s^{(1)}_{l}(k)/E[s^{(1)}_{1}(k)]$ with support on $l$ sites for the Heisenberg model with disorder $D=0,5,12.5,25$. The data for $D=5$ and $D=12.5$ are averaged over $50$, the data for $D=25$ over $70$ disorder realizations. There is no indication for a convergence with increasing order $k$ to an exponential decay with $l$ even for very strong disorder. For $D=0$ and $D=25$ the results for the non-interacting case are shown for comparison (symbols with dotted lines).
Whereas in the non-interacting, non-disordered case the norm of the terms supported on $l$ sites grows asymptotically in the same way independent of $l$ if the order of the commutator $k$ is much larger than $l$, see Eq. (III.12), $s^{(1)}_{l}(k)/s^{(1)}_{1}(k)$ has a maximum which shifts to larger and larger $l$ with increasing $k$ in the interacting case. For large $k$, the total norm is thus completely dominated by terms which have support on a large number of lattice sites. This is to be expected, given that the number of distinct terms in the commutator is increasing exponentially which also leads to a rapidly growing number of possible paths between those terms, see also the graph in Fig. 13. Note that the graph is no longer regular; the number of edges per node is not constant and is increasing with the level the node is situated on. Note that this increase of the number of edges is a prerequisite for a faster than exponential growth of the norm.
@@
III.2.2 Many-body localization
@@
From the perspective of operator spreading, localization in a many-body system is thus an extremely challenging requirement. In a graph with an equal number of edges at each level as realized in a non-interacting model, see Fig. 4, it is easy to see—at least in the limit of large disorder—how the norm of terms with support on only a few sites can dominate the total norm. The exponential growth of edges in the interacting case (see Fig. 13), on the other hand, suggests that the total norm for large $k$ will be entirely dominated by terms with support on many lattice sites. I.e., a local operator is generically expected to delocalize. This aspect also makes it difficult to see how a perturbative construction of local integrals of motion can be arranged. This is a point which seems to have not been sufficiently appreciated in Refs. [47, 48] where such a construction was suggested. We will get back to this point in Sec. IV. The only way a local operator can avoid to spread through the entire system in the many-body case thus seems to be intricate interference effects between the exponentially many different paths in the graph, Fig. 13. While such fine-tuned destructive interference seems implausible as a generic mechanism for many-body localization we cannot, in general, show that such effects do not exist. We note, however, that for the transverse Ising chain with random fields—the model considered in the proof of MBL in Refs. [47, 48]—a gauge can be defined such that the commutator has a stoquastic-type property [37]. I.e., all terms in the commutator in this gauge have a positive amplitude. For this model, interference effects as a mechanism for localization are thus strictly excluded. For the Heisenberg model considered here, such a gauge does not exist. We will instead consider data from symbolic computations up to order $k=16$ and show that there is no evidence for destructive interference effects and a corresponding localization. We note that at order $k$, we are considering operators which are spread over $12k+$ 1 sites in the lattice. I.e., at order $k=16$ we are dealing effectively with a chain of $33$ lattice sites which is well beyond what is possible in exact diagonalizations. We will also consider disorder strengths up to $D=25$ which is much larger than the critical values which have been suggested in the past [8, 10, 11, 49].
@@
Let us start by considering the number of distinct terms in the commutator at order $k$ which is shown in Fig. 9. We find that the number of terms grows faster in the disordered than in the case without disorder. This is easy to understand: additional terms appear in the commutator due to the random magnetic field terms while all the terms present before remain. The different classes of terms which appear up to order $k=3$ are shown in the graph in Fig. 13. Next, we consider the growth of the total norm which should be at most exponential in a localized phase as we have proven in Eq. (II.21), see also App. A. The data, however, are overall much better described by the general bound (II.12) than by an exponential, with the norm bound of the local Hamiltonian, $||h^{a}_{I}||\leq J$, used as a fitting parameter. While this does not exclude the possibility that for larger $k$ the scaling eventually does become exponential, we can say with confidence that no indications for localization are observed in spin chains up to $33$ sites. These system sizes are much larger than those used in exact diagonalizations, the method which claims of a many-body localized phase have been mostly based on [8, 10, 11, 19]. The stronger than exponential growth of the norm is further supported by considering the ratio of norms $s^{(2)}(k)/s^{(2)}(k-1)$ which is shown in Fig. 11. For the orders of the commutator which we can handle symbolically, there is no indication that this norm ratio saturates, even for extremely strong disorder, as it has to in a localized case where $s(k)\sim\exp(k)$. As the figure shows, the behavior in the interacting case is in stark contrast to the localized non-interacting case.
@@
Finally, let us also consider the norm ratio $s^{(1)}_{l}(k)/E[s^{(1)}_{1}(k)]$ of terms with support on $l$ sites which is shown in Fig. 12 for various disorder strengths. Here $E[s^{(1)}_{1}(k)]$ denotes the mean of terms with support on one site. Even for very strong disorder $D=25$ there is no indication of a convergence to an exponential decay with $l$. To the contrary, there is always a maximum which shifts to larger $l$ with increasing order $k$ just as in the case without disorder.
@@
To further highlight this point, we show in Fig. 14 the norm ratio $s^{(1)}_{l}(k)/s^{(1)}_{3}(k)$ as a function of the order $k$ for various fixed $l$. In contrast to the non-interacting case shown in Fig. 7, the curves do not saturate for large $k$ but rather are consistent with the weight in the total norm shifting to larger and larger $l$ with increasing $k$. While we cannot exclude that this trend changes for even larger $k$ than the ones we can handle in symbolic calculations, we want to stress once more that the terms in the commutator are spread over $33$ sites for $k=16$. This is already much larger than the system sizes typically considered in exact diagonalizations. At the very least, we can say that for chains of lengths up to $33$ sites, there is no indication that a many-body localized phase will be realized even at disorder strengths up to $D=25$.
IV Schrieffer-Wolff transformations
@@
Figure 15: Top left: Number of terms in the Hamiltonian with support on $l$ sites after performing the Schrieffer-Wolff transformation (IV.9) up to order $k$ for the XXX chain with disorder and $N=100$ lattice sites. For large $k$, this number increases approximately exponentially with $l$. Note that $\tilde{n}_{l}(k)$ does not depend on the disorder realization; the results are exact. Bottom left: Total norm of terms with support on one site after $k=3$ transformations for $200$ samples versus $\text{max}_{i}|h_{i}-h_{i+1}|^{-1}$. Right: Total norm of terms with support on $l$ sites after $k$ transformations averaged over $200$ realizations for $N=100$ and $D=5$. While each term is expected to have an amplitude $\sim D^{-2l}$ if there are no resonances, the exponentially increasing number of terms means that the total coupling of distant sites is no longer controlled by a small parameter. For example, we have $\tilde{s}_{1}^{(1)}(k)\approx\tilde{s}_{8}^{(1)}(k)$ for $k=4$.
@@
We have so far considered the commutator of a microscopic lattice Hamiltonian $H$ with a local operator $\sigma^{z}_{i}$. If this Hamiltonian is in a localized phase, then it is unitarily equivalent to a Hamiltonian of (interacting) conserved charges $\tau^{z}_{j}$ which are localized around lattice site $j$, see Eq. (I.2). Our argument has been that the local operator $\sigma^{z}_{i}$ then only has significant overlap with conserved charges $\tau^{z}_{j}$ which are localized near site $i$. Thus, the operator $\sigma^{z}_{i}$ will remain localized—up to exponential tails—under Euclidean time evolution.
@@
An alternative but related approach is to try to explicitly construct the unitarily equivalent Hamiltonian written in terms of local conserved charges $\tau^{z}_{j}$ [47, 48, 50, 51, 52, 29]. It is important to note that these charges are not uniquely defined: if, for example, $\tau^{z}_{j}$ and $\tau^{z}_{j+1}$ are conserved charges localized near lattice site $j$ with $[H,\tau^{z}_{j}]=[H,\tau^{z}_{j+1}]=0$ then $\tau^{z}_{j}\tau^{z}_{j+1}$ is also local and conserved, $[H,\tau^{z}_{j}\tau^{z}_{j+1}]=0$.
@@
$H\to\exp(T)H\exp(-T)=\sum_{k=0}^{\infty}\frac{[T,H]^{(k)}}{k!}$ (IV.1)
with $T^{\dagger}=-T$ being an anti-unitary operator. For a Hamiltonian $H=H_{0}+V$ where $H_{0}$ is diagonal in the operator basis chosen, we can instead of a full diagonalization demand that the new Hamiltonian $H^{(1)}$ is diagonal only to first order in the perturbation $V$. This is the Schrieffer-Wolff transformation [53, 54]. Going one step further, we can consider successive Schrieffer-Wolff transformations where in each step the perturbation is eliminated to leading order. This is closely related to the approach in Ref. [47, 48] which tries to generalize a proof of Anderson localization to the many-body case. In these works, the author is mostly concerned with the impact of resonances and splits the off-diagonal part of the Hamiltonian into a resonant and a non-resonant part. We will discuss the impact of resonances as well but will point out an even more fundamental issue which calls into question the entire perturbative construction of local conserved charges in the interacting case even for disorder configurations where no resonances are present.
To be specific, we consider again the XXZ Hamiltonian (III.1). We note that in Refs. [47, 48] the transverse field Ising model with random couplings and random fields was considered instead. However, the arguments made in the following are quite general and will apply to the latter model as well [37]. We want to investigate the case of large disorder where the magnetic field and the interaction terms constitute $H_{0}$ with $[H_{0},\sigma^{z}_{i}]=0$ and the hopping terms are the perturbation. The condition the operator $T$ has to fulfill to lowest order is then given by
$V+[T,H_{0}]=0$ (IV.2)
@@
$H^{(1)}=H_{0}+\frac{1}{2}[T,V]+\frac{1}{3}[T,[T,V]]+\cdots\,.$ (IV.3)
In the Schrieffer-Wolff transformation, one typically stops at the first order correction to $H_{0}$. To justify this perturbative approach, we can rescale the Hamiltonian (III.1) by $1/D$ which changes the energy scales but leaves the physical properties unchanged. Then, the hopping terms have an amplitude $\bar{J}^{(0)}_{j,j+1}=1/D$ and the interaction terms an amplitude $\bar{\Delta}=\Delta/D$ which are both small parameters for large $D$. After this rescaling, the random magnetic fields are always given by $h_{j}\in[-1,1]$.
IV.1 Non-interacting case
First, we consider again the non-interacting case. An ansatz for the operator $T$ can be obtained by commuting $H_{0}$ and $V$. This leads to $T=\sum_{j}A_{j}(\sigma^{+}_{j}\sigma^{-}_{j+1}-\sigma^{+}_{j+1}\sigma^{-}_{j})$. The amplitudes $A_{j}$ can then be determined from Eq. (IV.2) leading to
$T=\frac{1}{2}\sum_{j}\frac{\bar{J}^{(0)}_{j,j+1}}{h_{j}-h_{j+1}}(\sigma^{+}_{j}\sigma^{-}_{j+1}-\sigma^{+}_{j+1}\sigma^{-}_{j})\,.$ (IV.4)

### v4

@@ V.2.1 No disorder
From the exact solution of the model, we know for sure that the eigenstates of the XXX model are not localized. We can calculate the number of distinct terms $n(k)$ in the commutator $[H,\sigma_{0}^{z}]^{(k)}$ at order $k$ by symbolic calculations up to order $k=18$ exactly. We have not been able to identify the series but asymptotically the number of terms $n(k)$ does grow exponentially with a multiplicative algebraic correction which is consistent with the general bound (III.15) which we derived earlier. The data and a corresponding fit are shown in Fig. 12.
@@
Figure 12: In the interacting model with $\Delta=1$, the number of different terms $n(k)$ in the commutator $[H,\sigma_{0}^{z}]^{(k)}$ grows exponentially (symbols) and is very well fitted by $n(k)\sim 1.514\,k^{0.803}\exp(0.922k)$ in the non-disordered case and $n(k)\sim 1.492\,k^{0.815}\exp(0.998k)$ in the case with disorder (solid lines). The general bounds (III.15) with $f=3$ and $f=4$ are shown as well.
@@
The fit works extremely well with a least square deviation of $\chi^{2}\sim 10^{-2}$. The local operator basis for the Heisenberg model has dimension $f=4$ and consists of the operators ${\mathbbm{1},\sigma^{x},\sigma^{y},\sigma^{z}}$. We note, however, that only a small subset of terms in $[H,\sigma^{z}_{0}]^{(k)}$ contains an identity in between Pauli matrices. Ignoring such terms, we can set $f=3$ in Eq. (III.15) which gives a much tighter bound, see Fig. 12. The exponential growth of $n(k)$, instead of the quadratic growth in the free-fermion case, indicates that in the corresponding commutator graph the number of nodes and edges per level are no longer constant.
@@
We note that the number of distinct terms is different and always smaller than the number of connected clusters which we have calculated in Eq. (III.11). The reason is that different connected clusters can result in the same term. For the calculation of the norm this does not matter though. We can bound the contribution of each connected cluster as in Eq. (III.12) even if some of the resulting terms are the same. In Fig. 13 we show that this general bound indeed describes the norm growth well if we use the bound on the local Hamiltonian, $||h^{a}_{I}||\leq J$, as a fitting parameter.
@@
Figure 13: The 1-norm in the interacting case with $\Delta=1$ grows faster than exponential for all disorder strengths shown, consistent with the rescaled general bound, Eq. (III.12). The data are averaged over $30$ disorder realizations for $D=2.5$, $D=5$ and $D=12.5$ and over 50 realizations for $D=25$.
@@
Figure 15: The norm of operators $s^{(1)}_{l}(k)/E[s^{(1)}_{1}(k)]$ with support on $l$ sites for the Heisenberg model with disorder $D=0,5,12.5,25$. The data for $D=5$ and $D=12.5$ are averaged over $50$, the data for $D=25$ over $70$ disorder realizations. There is no indication for a convergence with increasing order $k$ to an exponential or almost exponential decay with $l$ even for very strong disorder. For $D=0$ and $D=25$ the results for the non-interacting case are shown for comparison (symbols with dotted lines).
Whereas in the non-interacting, non-disordered case the norm of the terms supported on $l$ sites grows asymptotically in the same way independent of $l$ if the order of the commutator $k$ is much larger than $l$, see Eq. (V.12), $s^{(1)}_{l}(k)/s^{(1)}_{1}(k)$ has a maximum which shifts to larger and larger $l$ with increasing $k$ in the interacting case. For large $k$, the total norm is thus completely dominated by terms which have support on a large number of lattice sites. This is to be expected, given that the number of distinct terms in the commutator is increasing exponentially which also leads to a rapidly growing number of possible paths between those terms, see also the graph in Fig. 16. Note that the graph is no longer regular; the number of edges per node is not constant and is increasing with the level the node is situated on. Note that this increase of the number of edges is a prerequisite for a faster than exponential growth of the norm.
@@
V.2.2 Many-body localization
@@
Based on this graph representation, it is obvious that a strict exponential localization of operators in a many-body system is thus an extremely challenging requirement. In a graph with an equal number of edges at each level as realized in a non-interacting model, see Fig. 7, it is easy to see—at least in the limit of large disorder—how the norm of terms with support on only a few sites can dominate the total norm. The exponential growth of edges in the interacting case (see Fig. 16), on the other hand, suggests that the total norm for large $k$ will be entirely dominated by terms with support on many lattice sites. I.e., a local operator is generically expected to delocalize. This aspect also makes it difficult to see how a perturbative construction of local integrals of motion can be arranged. This is a point which seems to have not been sufficiently appreciated in Refs. [48, 49] where such a construction was suggested. We will get back to this point in Sec. VI. The only way a local operator can avoid to spread through the entire system in the many-body case thus seems to be intricate interference effects between the exponentially many different paths in the graph, Fig. 16. While such fine-tuned destructive interference seems implausible as a generic mechanism for many-body localization we cannot, in general, show that such effects do not exist. We note, however, that for the transverse Ising chain with random fields—the model considered in the proof of MBL in Refs. [48, 49]—a gauge can be defined such that the commutator has a stoquastic-type property [37]. I.e., all terms in the commutator in this gauge have a positive amplitude. For this model, interference effects as a mechanism for localization are thus strictly excluded and we have shown in Sec. IV that the almost factorial norm growth is indeed incompatible with exponential localization. For the Heisenberg model considered here, such a gauge does not exist. We will instead consider data from symbolic computations up to order $k=16$ and show that there is no evidence for destructive interference effects and a corresponding localization. The data are consistent with a stronger than exponential norm growth. We note that at order $k$, we are considering operators which are spread over $12k+$ 1 sites in the lattice. I.e., at order $k=16$ we are dealing effectively with a chain of $33$ lattice sites which is well beyond what is possible in exact diagonalizations. We will also consider disorder strengths up to $D=25$ which is much larger than the critical values which have been suggested in the past [8, 10, 11, 50].
@@
Let us start by considering the number of distinct terms in the commutator at order $k$ which is shown in Fig. 12. We find that the number of terms grows faster in the disordered than in the case without disorder. This is easy to understand: additional terms appear in the commutator due to the random magnetic field terms while all the terms present before remain. The different classes of terms which appear up to order $k=3$ are shown in the graph in Fig. 16. Next, we consider the growth of the total norm which should be at most exponential in a strictly exponentially localized phase as we have proven in Eq. (IV.4). The data, however, are overall much better described by the general bound (III.12) than by an exponential, with the norm bound of the local Hamiltonian, $||h^{a}_{I}||\leq J$, used as a fitting parameter. While this does not exclude the possibility that for larger $k$ the scaling eventually does become exponential, we can say with confidence that no indications for a strict exponential localization are observed in spin chains up to $33$ sites. These system sizes are much larger than those used in exact diagonalizations, the method which claims of a many-body localized phase have been mostly based on [8, 10, 11, 19]. The stronger than exponential growth of the norm is further supported by considering the ratio of norms $s^{(2)}(k)/s^{(2)}(k-1)$ which is shown in Fig. 14. For the orders of the commutator which we can handle symbolically, there is no indication that this norm ratio saturates, even for extremely strong disorder, as it has to in the strictly localized case where $s(k)\sim\exp(k)$. As the figure shows, the behavior in the interacting case is in stark contrast to the localized non-interacting case.
@@
Finally, let us also consider the norm ratio $s^{(1)}_{l}(k)/E[s^{(1)}_{1}(k)]$ of terms with support on $l$ sites which is shown in Fig. 15 for various disorder strengths. Here $E[s^{(1)}_{1}(k)]$ denotes the mean of terms with support on one site. Even for very strong disorder $D=25$ there is no indication of a convergence to an exponential decay with $l$ as is expected in the strictly exponentially localized case. To the contrary, there is always a maximum which shifts to larger $l$ with increasing order $k$ just as in the case without disorder.
@@
To further highlight this point, we show in Fig. 17 the norm ratio $s^{(1)}_{l}(k)/s^{(1)}_{3}(k)$ as a function of the order $k$ for various fixed $l$. In contrast to the non-interacting case shown in Fig. 10, the curves do not saturate for large $k$ but rather are consistent with the weight in the total norm shifting to larger and larger $l$ with increasing $k$. While we cannot exclude that this trend changes for even larger $k$ than the ones we can handle in symbolic calculations, we want to stress once more that the terms in the commutator are spread over $33$ sites for $k=16$. This is already much larger than the system sizes typically considered in exact diagonalizations. At the very least, we can say that for chains of lengths up to $33$ sites, there is no indication that a phase with strict exponential localization is realized even at disorder strengths up to $D=25$.
This leaves open the second scenario where the Hamiltonian is unitarily equivalent to the effective Hamiltonian $\tilde{H}$ and where instead of a strict exponential localization, local operators spread across the lattice according to Eq. (II.4). First, we note that the decay rate $\tilde{\kappa}_{2}=\tilde{\kappa}_{2}(D)$ is increasing with increasing $D$ so that for very strong disorder there should be an almost exponential decay with $l$. There is no indication for such a behavior in Fig. 15 even for $D=25$. However, it is of course possible that $D$ is still not large enough. Second, we note that Eq. (IV.2) predicts a maximum at $l=k/\tilde{\kappa}_{2}$. This is not inconsistent with Fig. (15) where we indeed observe an approximately linear shift of the maximum with increasing commutator order $k$. For the norm ratios shown in Fig. 17 the result (IV.2) predicts that $s_{l}(k)/s_{3}(k)\sim l^{k}$ which is also not inconsistent with the data from the symbolic calculations (note the logarithmic scale).
To summarize, the symbolic data are consistent with a maximal total norm growth in the disordered Heisenberg chain for all disorder strengths as in the random field Ising model. This and the data for the norms with support on $l$ sites seem to exclude the existence of a phase where operators remain strictly exponentially localized. However, the data are not inconsistent with the second scenario where $H$ is unitarily equivalent to the effective Hamiltonian $\tilde{H}$ describing a non-ergodic phase where local operators, however, eventually spread over the entire lattice and do not remain localized. We therefore investigate next if the unitary transform $U$ can be constructed perturbatively.
VI Schrieffer-Wolff transformations
@@
Figure 18: Top left: Number of terms in the Hamiltonian with support on $l$ sites after performing the Schrieffer-Wolff transformation (VI.9) up to order $k$ for the XXX chain with disorder and $N=100$ lattice sites. For large $k$, this number increases approximately exponentially with $l$. Note that $\tilde{n}_{l}(k)$ does not depend on the disorder realization; the results are exact. Bottom left: Total norm of terms with support on one site after $k=3$ transformations for $200$ samples versus $\text{max}_{i}|h_{i}-h_{i+1}|^{-1}$. Right: Total norm of terms with support on $l$ sites after $k$ transformations averaged over $200$ realizations for $N=100$ and $D=5$. While each term is expected to have an amplitude $\sim D^{-2l}$ if there are no resonances, the exponentially increasing number of terms means that the total coupling of distant sites is no longer controlled by a small parameter. For example, we have $\tilde{s}_{1}^{(1)}(k)\approx\tilde{s}_{8}^{(1)}(k)$ for $k=4$.
@@
We have so far considered the commutator of a microscopic lattice Hamiltonian $H$ with a local operator. If this Hamiltonian is in a strictly exponentially localized phase, then the operator will remain localized—up to exponential tails—under Euclidean time evolution. However, we have also seen that if we only demand that the Hamiltonian is unitarily equivalent to an effective Hamiltonian of (interacting) conserved charges, then this might open up the possibility for a non-ergodic phase where local operators, however, do not remain localized. Our symbolic calculations of the commutator were unable to completely rule out this scenario.
@@
We therefore consider next if we can explicitly construct the unitarily equivalent Hamiltonian $\tilde{H}$ written in terms of local conserved charges $\tau^{z}_{j}$ [48, 49, 51, 52, 53, 29]. It is important to note that these charges are not uniquely defined: if, for example, $\tau^{z}_{j}$ and $\tau^{z}_{j+1}$ are conserved charges localized near lattice site $j$ with $[H,\tau^{z}_{j}]=[H,\tau^{z}_{j+1}]=0$ then $\tau^{z}_{j}\tau^{z}_{j+1}$ is also local and conserved, $[H,\tau^{z}_{j}\tau^{z}_{j+1}]=0$.
@@
$H\to\exp(T)H\exp(-T)=\sum_{k=0}^{\infty}\frac{[T,H]^{(k)}}{k!}$ (VI.1)
with $T^{\dagger}=-T$ being an anti-unitary operator. For a Hamiltonian $H=H_{0}+V$ where $H_{0}$ is diagonal in the operator basis chosen, we can instead of a full diagonalization demand that the new Hamiltonian $H^{(1)}$ is diagonal only to first order in the perturbation $V$. This is the Schrieffer-Wolff transformation [54, 55]. Going one step further, we can consider successive Schrieffer-Wolff transformations where in each step the perturbation is eliminated to leading order. This is closely related to the approach in Ref. [48, 49] which tries to generalize a proof of Anderson localization to the many-body case. In these works, the author is mostly concerned with the impact of resonances and splits the off-diagonal part of the Hamiltonian into a resonant and a non-resonant part. We will discuss the impact of resonances as well but will point out an even more fundamental issue which might call into question the entire perturbative construction of local conserved charges in the interacting case even for disorder configurations where no resonances are present.
To be specific, we consider again the XXZ Hamiltonian (V.1). We note that in Refs. [48, 49] the transverse field Ising model with random couplings and random fields was considered instead. However, the arguments made in the following are quite general and will apply to the latter model as well [37]. We want to investigate the case of large disorder where the magnetic field and the interaction terms constitute $H_{0}$ with $[H_{0},\sigma^{z}_{i}]=0$ and the hopping terms are the perturbation. The condition the operator $T$ has to fulfill to lowest order is then given by
$V+[T,H_{0}]=0$ (VI.2)
@@
$H^{(1)}=H_{0}+\frac{1}{2}[T,V]+\frac{1}{3}[T,[T,V]]+\cdots\,.$ (VI.3)
In the Schrieffer-Wolff transformation, one typically stops at the first order correction to $H_{0}$. To justify this perturbative approach, we can rescale the Hamiltonian (V.1) by $1/D$ which changes the energy scales but leaves the physical properties unchanged. Then, the hopping terms have an amplitude $\bar{J}^{(0)}_{j,j+1}=1/D$ and the interaction terms an amplitude $\bar{\Delta}=\Delta/D$ which are both small parameters for large $D$. After this rescaling, the random magnetic fields are always given by $h_{j}\in[-1,1]$.
VI.1 Non-interacting case
First, we consider again the non-interacting case. An ansatz for the operator $T$ can be obtained by commuting $H_{0}$ and $V$. This leads to $T=\sum_{j}A_{j}(\sigma^{+}_{j}\sigma^{-}_{j+1}-\sigma^{+}_{j+1}\sigma^{-}_{j})$. The amplitudes $A_{j}$ can then be determined from Eq. (VI.2) leading to
$T=\frac{1}{2}\sum_{j}\frac{\bar{J}^{(0)}_{j,j+1}}{h_{j}-h_{j+1}}(\sigma^{+}_{j}\sigma^{-}_{j+1}-\sigma^{+}_{j+1}\sigma^{-}_{j})\,.$ (VI.4)